### v1

# Nonlinear semiclassical spectroscopy of ultrafast molecular polariton dynamics

Abstract
We introduce a theoretical framework that allows for the systematic and efficient description of the ultrafast nonlinear response of molecular polaritons, i.e., hybrid light-matter states, in the collective regime of large numbers of molecules $\mathcal{N}$ coupled to the cavity photon mode. Our approach is based on a semiclassical, mean-field evolution of the molecular Hamiltonian and the cavity field, which is complemented by a perturbative expansion of both light and matter counterparts in the input fields entering the cavity. In addition, expansion in terms of the pulse phases enables us to disentangle different excitation pathways in Liouville space, thereby distinguishing contributions to the nonlinear response. The formalism extends traditional free-space nonlinear spectroscopy by incorporating the feedback of matter onto the light field via the induced polarization. We demonstrate the utility of the framework by applying it to the calculation of pump-probe polariton spectra and show how, by storing the pulses, the cavity facilitates additional excitation pathways which are not possible in free space. Our method, which does not scale with $\mathcal{N}$, is broadly applicable and can be extended to a wide range of current experiments investigating the dynamical nonlinear response of hybrid light-matter states.
pacs:
42.50.Pq, 42.65.-k, 42.65.Sf
Figure 1: (a) Schematics. An optical cavity containing an ensemble of $\mathcal{N}$ molecules is driven by input pulses with amplitudes $\eta_{p}$ (pump) and $\eta_{p^{\prime}}$ (probe). (b) Key distinction between conventional, free-space nonlinear spectroscopy and nonlinear cavity spectroscopy, concerning the evolution of the molecular density matrix. For cavity spectroscopy, the feedback of the molecular polarization onto the cavity field leads to a coupling of the $n$-th order molecular density matrix $\vec{\rho}^{(n)}$ to all lower orders in the cavity field, as opposed to the free-space case, where the interaction is always linear in the laser field that drives the molecules. (c) Pump and probe pulses entering the cavity with time delay $\tau_{\Delta}$ get stored and prolonged in the cavity on a time scale given by the cavity lifetime $\kappa^{-1}$, therefore allowing the cavity field created by the probe ($\alpha^{(0)(1)}$) to act before the pump ($\alpha^{(1)(0)}$). Additional phases $\Phi_{p,p^{\prime}}$ imprinted on the pulses can be utilized to separate excitation pathways in Liouville space.
Introduction.—Molecular polaritons, quasiparticles resulting from the strong coupling between confined photonic modes and vibrational or electronic excitations of molecules, e.g., inside Fabry-Pérot cavities, have gained significant attention over recent years, largely due to their potential applications in areas such as polariton chemistry Hutchison et al. (2012); Thomas et al. (2019), energy transport Coles et al. (2014); Zhong et al. (2017), condensation and lasing Kéna-Cohen and Forrest (2010); Plumhof et al. (2014), and the manipulation of optical nonlinearities Barachati et al. (2018); Xiang et al. (2019); Wang et al. (2021); Cheng et al. (2022). Recent advances in electronic Takemura et al. (2015a, b); DelPo et al. (2020); Fassioli et al. (2021); Mewes et al. (2020); Son et al. (2022); Wu et al. (2022); Russo et al. (2024) and vibrational Xiang et al. (2018, 2019); Xiong (2023); Sufrin et al. (2024); Dunkelberger et al. (2022); Duan et al. (2021); Simpkins et al. (2023) nonlinear polariton spectroscopy have provided new insight into the nonlinear response of molecular polaritons by revealing the dynamics of the complex interactions and energy relaxation pathways. However, significant challenges and open questions remain in understanding the nonlinear polariton response, particularly regarding the short-time dynamical behavior, i.e., before the cavity photons, and thus the polariton states, have decayed ($\sim$ fs in experiments). The development of an efficient formalism that can address these questions is crucial for advancing the understanding of polariton dynamics across different time scales as observed in pump-probe experiments.
While the theoretical formalism of nonlinear spectroscopy involving the interaction between laser light and matter in free space is well-established Mukamel (1995); Yuen-Zhou et al. (2014); Jonas (2003); Cho (2008); Gelin et al. (2009), recent works have aimed at constructing a framework for the theory of of nonlinear electronic and vibrational polariton spectroscopy Saurabh and Mukamel (2016); F. Ribeiro et al. (2018); Zhang et al. (2023); Mondal et al. (2023); Gallego-Valencia et al. (2024); Schnappinger et al. (2024). Some studies have considered direct excitation of the molecules through the side of the cavity Zhang et al. (2023), whereas other works have taken a fully quantum mechanical approach, thereby restricting the scope of the treatment to either long times after light and matter have decoupled, or the explicit simulation of only a few molecules F. Ribeiro et al. (2018); Gallego-Valencia et al. (2024).
In this Letter, we introduce a systematic formalism to nonlinear cavity spectroscopy in the limit of large numbers of molecules $\mathcal{N}$ per photon mode. Our approach is based on a mean-field evolution of the coupled light-matter system Fowler-Wright et al. (2022), related to semiclassical Maxwell-Liouville approaches which have been developed, e.g., in semiconductor optics and nanophotonics Jahnke et al. (1996); Scully and Zubairy (1997); Lopata and Neuhauser (2009); Šindelka (2010); Sukharev and Nitzan (2011); Li et al. (2018); Chen et al. (2019). We perturbatively expand both cavity field and density matrix in the input fields driving the cavity and show that the formalism can be regarded as an extension of the conventional free-space perturbative nonlinear spectroscopy framework by Mukamel Mukamel (1995), now incorporating nonlinear contributions to the field resulting from the feedback of the molecules onto the cavity dynamics. In addition, expansion of the nonlinear response in terms of phase components allows us to separate different excitation pathways in Liouville space and identify new pathways enabled by the cavity, using the widely-adopted toolkit of double-sided Feynman diagrams.
Mean-field approach.—We consider an ensemble of $\mathcal{N}$ molecules collectively coupled to the single confined mode of an optical cavity at frequency $\omega_{c}$ [see schematics in Fig. 1(a)]. Each molecule $j$ is described by an (in principle arbitrary) Hamiltonian $\mathcal{H}_{0}^{j}$ with a corresponding dipole operator $\hat{\mu}_{j}$. The interaction between molecules and the cavity field is described by a Dicke-type Hamiltonian such that the total Hamiltonian describing the full quantum evolution of the coupled light-matter system expresses as (we set $\hbar\equiv 1$)
$\mathcal{H}=\omega_{c}\hat{a}^{\dagger}\hat{a}+\sum_{j=1}^{\mathcal{N}}\mathcal{H}_{0}^{j}+E_{0}(\hat{a}+\hat{a}^{\dagger})\sum_{j=1}^{\mathcal{N}}\hat{\mu}_{j},$ (1)
with the cavity zero-point amplitude $E_{0}=\sqrt{\omega_{c}/(2\epsilon_{0}\mathcal{V})}$, where $\mathcal{V}$ is the cavity mode volume and $\epsilon_{0}$ is the vacuum permittivity. To facilitate the understanding of the formalism, we initially assume that the cavity field is driven by only a single input field with amplitude $\eta$ at a central carrier frequency $\omega_{\ell}$ and with a Gaussian pulse shape described by the temporal envelope function $f(t)$ with a temporal pulse width $\tau_{w}$.
The limit of large $\mathcal{N}$ allows us to apply a mean-field approximation (exact for $\mathcal{N}\to\infty$), which amounts to setting $\expectationvalue{\hat{a}}=\alpha(t)$ Fowler-Wright et al. (2022). Additionally, for illustration, we take all molecules to be identical, i.e., $\hat{\mu}_{j}=\hat{\mu}$, although extension to disordered ensembles offers no conceptual difficulty. The simplified mean-field Hamiltonian then expresses as
$\mathcal{H}_{\text{MF}}=\mathcal{H}_{0}+E_{0}\left(\alpha(t)+\alpha^{*}(t)\right)\hat{\mu}.$ (2)
The dynamical evolution of the (now single-particle) molecular density matrix $\rho$ is then governed by the master equation $\partial_{t}\rho=\mathrm{i}[\rho,\mathcal{H}_{\text{MF}}]+\mathcal{D}[\rho]$. Here, the dissipator $\mathcal{D}[\rho]$ represents possible additional loss channels of the molecule (e.g., dephasing at rate $\gamma_{\phi}$, non-radiative decay,$\ldots$) that may be added, e.g., in Lindblad form. However, a complete description of the mean-field problem also involves solving the classical equation for the cavity field amplitude which undergoes damping at rate $\kappa$ (photon loss)
$\dot{\alpha}(t)$ $=-\left(\frac{\kappa}{2}+\mathrm{i}\omega_{c}\right)\alpha(t)-\mathrm{i}\mathcal{N}E_{0}P(t)-\eta f(t)\mathrm{e}^{-\mathrm{i}\omega_{\ell}t},$ (3)
containing the induced molecular polarization $P(t)=\mathrm{Tr}[\hat{\mu}\rho(t)]$ created by all $\mathcal{N}$ molecules. Altogether, this describes a self-consistent, semiclassical evolution of the coupled light-matter system, where the matter component evolves quantum mechanically, while the part describing the electromagnetic field undergoes classical evolution and is driven by both the input field as well as by the collective molecular polarization feeding back into the cavity field.
Perturbative expansion.—While the above set of mean-field equations can be readily solved numerically, there is generally no closed analytical solution due to the nonlinearity of the molecular transitions. To gain more insight into the structure of the nonlinear response, we therefore proceed in analogy to standard nonlinear spectroscopy with a perturbative expansion of both the field and the matter density matrix in terms of the input field amplitude, i.e.,
$\alpha(t)=\sum_{n=0}^{\infty}\eta^{n}\alpha^{(n)}(t),\quad\rho(t)=\sum_{n=0}^{\infty}\eta^{n}\rho^{(n)}(t).$ (4)
From this, also other perturbative quantities can be derived such as, e.g., the $n$-th order polarization $P^{(n)}(t)=\mathrm{Tr}[\hat{\mu}\rho^{(n)}(t)]$  [^1] Note that the $\rho^{(n)}$ are not true density matrices, i.e., they do not obey the properties of density matrices such as the conservation of the trace.
Furthermore, for both numerical and analytical considerations, it is advantageous to convert the master equation for the density matrix onto matrix-vector form, i.e., map the density matrix $\rho$ onto a vector $\vec{\rho}$ and the superoperator acting on the density matrix onto a matrix (also referred to as Liouville space) [see Supplementary Material (SM) sm for details on vectorization] Mukamel (1995); Horn and Johnson (1994); Am-Shallem et al. (2015). Then, one obtains a closed system for the cavity field amplitude and the vectorized density matrix up to $n$-th order
$\dot{\alpha}^{(n)}$ $=-\left(\frac{\kappa}{2}+\mathrm{i}\omega_{c}\right)\alpha^{(n)}-\mathrm{i}\mathcal{N}E_{0}P^{(n)}-\delta_{1n}f(t)\mathrm{e}^{-\mathrm{i}\omega_{\ell}t},$ (5a)
$\dot{\vec{\rho}}^{(n)}$ $=-\mathrm{i}\mathcal{L}_{0}\vec{\rho}^{(n)}-\mathrm{i}E_{0}\sum_{j=0}^{n}\left(\alpha^{(n-j)}+\alpha^{(n-j)*}\right)\mathcal{L}_{\mu}\vec{\rho}^{(j)},$ (5b)
in terms of the vectorized commutators $[\mathcal{H}_{0},\rho^{(n)}]\to\mathcal{L}_{0}\vec{\rho}^{(n)}$ and $[\hat{\mu},\rho^{(j)}]\to\mathcal{L}_{\mu}\vec{\rho}^{(j)}$ where $\mathcal{L}_{0}$, $\mathcal{L}_{\mu}$ are the free molecular and interaction Liouvillians, respectively. In particular, Eqs. (5) indicate that only the first-order cavity field is driven directly by the input fields while for higher orders, the molecular density matrix is driven by lower orders of the cavity field. Importantly, Eq. (5b) shows that the (cavity) field driving the molecular density matrix can have nonlinear contributions in the input field, as opposed to free-space spectroscopy where the interaction Liouvillian driving the evolution of the density matrix is always linear in the (laser) input field [see Fig. 1(b)] Mukamel (1995). In other words, the electric field experienced by the molecules is not the laser field interacting with the cavity from outside, but a distorted version of it due to the evolving material polarization inside the cavity. While for conventional spectroscopy with well-separated pulses, only a single multidimensional integral is obtained as solution for the $n$-th order density matrix, many additional integrals arise in the cavity scenario ($12^{n-1}-$ 1, see diagram in Fig. 2 and SM sm ). Also, a straightforward analysis of Eqs. (5) reveals that the first order ($n=1$) recovers the standard linear polariton response, characterized by a Rabi splitting proportional to $\sqrt{\mathcal{N}}$, arising from the hybridization of the cavity with the collective molecular bright mode Rocca et al. (1998); Agranovich et al. (2003); Litinskaya et al. (2004); Gonzalez-Ballestero et al. (2016); Campos-Gonzalez-Angulo and Yuen-Zhou (2022); Yuen-Zhou and Koner (2024).
Figure 2: Density matrix evolution. Graph illustrating the possible integrals contributing to the dynamical evolution of the $n$-th order density matrix $\rho^{(n)}(t)$. The orange path on the right illustrates the integral that is obtained for conventional free-space spectroscopy, corresponding to a sequence of first-order interactions linear in the input. The feedback of the nonlinear polarization onto the cavity field adds $12^{n-1}-$ 1 other possible pathways.
Figure 3: Differential transmission. DT spectra $\Delta T(\omega)$ for a 2LS with resonance frequency $\omega_{0}$ as a function of the delay time between the pump and probe pulse $\tau_{\Delta}$ for (a) weak ($\gamma_{\phi}=0$), and (b) strong ($\gamma_{\phi}=\kappa$) dephasing of the electronic transition. The green dashed lines show the location of the (linear response) polariton frequencies. The black curve in (b) shows a cross section through the DT at $\tau_{\Delta}=4\kappa^{-1}$. The plots in (c) and (d) show the dynamics of the population and polarization $|P(t)|^{2}$ created by the initial pump pulse arriving at $\tau_{p}=2\kappa^{-1}$ which are then probed by the second pulse after $\tau_{\Delta}$. The 2LS is initialized in the ground state, with parameters $g\sqrt{\mathcal{N}}=3\kappa$, $\tau_{w}=0.05\kappa^{-1}$, $\omega_{0}=\omega_{c}$, and central frequency $\omega_{p}=\omega_{p^{\prime}}=\omega_{0}$ for both pulses.
Differential transmission.—We now consider two pulses driving the cavity (pump $p$ and probe $p^{\prime}$), with input fields proportional to $\eta_{j}f_{j}(t-\tau_{j})\mathrm{e}^{-\mathrm{i}\omega_{j}t}$ where $\eta_{j}$, $f_{j}$, $\omega_{j}$, $\tau_{j}$ for $j\in\{p,p^{\prime}\}$ denote the amplitude, envelope, carrier frequency, and arrival time of the pump and probe pulses, respectively. We assume $\tau_{p^{\prime}}>\tau_{p}$, with a time delay between pump and probe of $\tau_{\Delta}=\tau_{p^{\prime}}-\tau_{p}$ [see sketch in Fig. 1(c)]. We proceed with a perturbative expansion in both pump ($p$) and probe ($p^{\prime}$) amplitudes (this can be generalized to arbitrary many pulses)
$\alpha(t)$ $=\sum_{n,m=0}^{\infty}\eta_{p}^{n}\eta_{p^{\prime}}^{m}\alpha^{(n)(m)}(t),$ (6a)
$\rho(t)$ $=\sum_{n,m=0}^{\infty}\eta_{p}^{n}\eta_{p^{\prime}}^{m}\rho^{(n)(m)}(t),$ (6b)
where the first and second indeces refer to the orders in the pump and probe fields, respectively. This allows us to derive a perturbative set of equations for $\alpha^{(n)(m)}$, $\vec{\rho}^{(n)(m)}$ similar to Eqs. (5), now describing pump-probe spectroscopy (see SM sm for details).
Figure 4: Separating phase contributions to the nonlinear polariton response. (a), (b) Plot of phase contributions $(0,1)$ and $(2,-1)$ to $\Delta T(\omega)$ as a function of the delay time between the pulses. The $(2,-1)$-contribution decays on a time scale of $\kappa^{-1}$, as it relies on the probe acting prior to the pump. The green dashed lines show the (linear response) polariton frequencies. We chose $g\sqrt{\mathcal{N}}=3\kappa$, $\gamma_{\phi}=0.1\kappa$, $\omega_{p}=\omega_{p^{\prime}}=\omega_{0}$, and a temporal width of $\tau_{w}=0.1\kappa^{-1}$ for both pulses. The double-sided Feynman diagrams in (c) and (d) represent processes contributing to the different phase combinations of $\alpha^{(2)(1)}$. The dotted arrows indicate the decay of the coherence.
From the solution for the total cavity field, the cavity transmission and reflection can be obtained using standard input-output theory Steck (2007). Since nonlinearities are typically weak, instead of absolute quantities, differential (nonlinear) corrections to the response are commonly considered. For the collinear geometry in question, we define the differential transmission (DT) spectrum by subtracting the (normalized) probe transmission without the pump from the probe transmission in the presence of the pump pulse
$\Delta T(\omega)=T_{p^{\prime}}^{\text{pump\, on}}(\omega)-T_{p^{\prime}}^{\text{pump\, off}}(\omega),$ (7)
where the transmission is related to the intracavity field $T_{p^{\prime}}=|(\kappa/2)\alpha(\omega)/[\eta_{p^{\prime}}f_{p^{\prime}}(\omega)]|^{2}$. The contribution $T_{p^{\prime}}^{\text{pump\, on}}(\omega)$ depends on the time delay between the pulses $\tau_{\Delta}$. By considering only the lowest (third) order correction due to pump and probe, the DT may be expressed as (see SM for derivation sm )
$\Delta T(\omega)\approx\left(\frac{\kappa}{2}\right)^{2}\eta_{p}^{2}\frac{2\mathrm{Re}\left[\alpha^{(0)(1)*}(\omega)\alpha^{(2)(1)}(\omega)\right]}{f_{p^{\prime}}(\omega)^{2}},$ (8)
where the key quantity is the third-order cavity field $\alpha^{(2)(1)}$. Beware that Eq. (8) is not the usual pump-probe signal; instead, it is a signal which contains different phase contributions (see paragraph below). The DT is plotted as a function of delay time in Figs. 3(a), (b) for two-level systems (2LSs) with transition dipole moment $\mu_{ge}$ coupled to the cavity at strength $g=\mu_{ge}E_{0}$, for various degrees of dephasing of the transition. The plots below in Figs. 3(c), (d) show the dynamics of the polarization and population created by an initial pump pulse which are then probed after a time $\tau_{\Delta}$. At early times, oscillatory behavior around the polariton frequencies is observed, corresponding to coherences between upper and lower polariton. Dephasing leads to a decay of coherence and the creation of stationary population in the molecules, even after the cavity photon has decayed. In the long-time limit, this manifests as a contracted Rabi splitting, characterized by a DT as shown by the black cross-section curve in Fig. 3(b) Xiang et al. (2018); Dunkelberger et al. (2022); F. Ribeiro et al. (2018).
Separating excitation pathways.—The nonlinear signal stemming from the $n$-th order density matrix generally contains many possible excitation pathways. For instance, even for free-space spectroscopy of a 2LS, there are eight Feynman diagrams contributing to the third-order nonlinear response Mukamel (1995). In experiments, it is therefore highly desired to isolate certain excitation pathways and thereby understand the different contributions to the nonlinear signal, which, owing to the collinear geometry in question, can be achieved by phase cycling. This method is equivalent to the more standard phase matching procedure, which, however obfuscates our formalism as it requires the description of a multimode cavity Yuen-Zhou et al. (2014); Jonas (2003); Hybl et al. (1998); Tian et al. (2003); Tan (2008); Yang et al. (2020).
To this end, we will consider the effect of additional phases applied to the pump ($\Phi_{p}$) and probe ($\Phi_{p^{\prime}}$) pulses. These phases are subsequently also imprinted onto the cavity field and the molecular density matrix, enabling us to disentangle different excitation pathways characterized by specific phase sequences Mukamel (1995); Yuen-Zhou et al. (2014). The cavity field and density matrix are now expanded as
$\alpha(t)$ $=\sum_{n,m,\vec{v}}\eta_{p}^{n}\eta_{p^{\prime}}^{m}\mathrm{e}^{-\mathrm{i}\vec{v}\cdot\vec{\Phi}}\alpha^{(n)(m)}_{\vec{v}}(t),$ (9a)
$\rho(t)$ $=\sum_{n,m,\vec{v}}\eta_{p}^{n}\eta_{p^{\prime}}^{m}\mathrm{e}^{-\mathrm{i}\vec{v}\cdot\vec{\Phi}}\rho^{(n)(m)}_{\vec{v}}(t),$ (9b)
where $\vec{\Phi}=(\Phi_{p},\Phi_{p^{\prime}})^{\top}$ denotes the vector of pump and probe phases, and $\vec{v}=(v_{p},v_{p^{\prime}})^{\top}$ with $v_{p},v_{p^{\prime}}\in\mathbb{Z}_{0}$ denotes the phase coefficients which also depend on the pump and probe orders $v_{p}(n)$, $v_{p^{\prime}}(m)$ (see SM sm ). The raw DT signal from Eq. (8) can then be separated into different phase contributions
$\Delta T_{\vec{v}}(\omega)\approx\left(\frac{\kappa}{2}\right)^{2}\eta_{p}^{2}\frac{2\mathrm{Re}\left[\alpha^{(0)(1)*}_{(0,-1)}(\omega)\alpha^{(2)(1)}_{\vec{v}}(\omega)\right]}{f_{p^{\prime}}(\omega)^{2}},$ (10)
with the total signal simply being obtained as the sum of all contributions $\Delta T=\sum_{\vec{v}}\Delta T_{\vec{v}}$. One finds that there are two possible phase combinations contributing to $\alpha^{(2)(1)}$: $(0,1)$, and $(2,-1)$ [and the complex conjugate processes $(0,-1)$, $(-2,1)$], which are shown in Figs. 4(a), (b), and are associated with pump-probe and double-quantum coherence (DQC) spectroscopies, respectively. The first contribution [$(0,1)\equiv\pm\Phi_{p}\mp\Phi_{p}+\Phi_{p^{\prime}}$] is the ‘standard’ pump-probe, where populations and zeroth-order quantum coherences are created by twofold application of the pump with opposite phase [see Feynman diagrams in Fig. 4(c)]. The coherent superposition between upper and lower polaritons then undergo Rabi oscillations which are detected by the probe. The second phase combination [($(2,-1)\equiv 2\Phi_{p}-\Phi_{p^{\prime}}$] is negligible outside of the cavity, as it implies that the 2LS is excited twice. However, inside the cavity, since the pulses get extended on a time scale of $\kappa^{-1}$ [cf. Fig. 1(c)], the probe can act before the pump, therefore enabling this phase combination to arise within $\kappa^{-1}$. In the SM sm , we derive the phase contributions up to third order for an arbitrary multi-level system and provide an illustration for a slightly more complicated three-level system, where the $(2,-1)$-contribution (DQC) can also exist outside of the cavity.
Discussion and conclusions.—We have presented an efficient and broadly applicable framework for nonlinear polariton spectroscopy, based on a mean-field evolution of the coupled light-matter system. The formalism extends conventional free-space spectroscopy by accounting for the feedback of the molecular polarization onto the cavity field and allows for an understanding of the polariton response beyond the stationary limit F. Ribeiro et al. (2018). Expanding the cavity field and molecular density matrix in terms of the input fields’ amplitudes and phases allows for a systematic understanding of the various contributions and excitation pathways to the nonlinear response.
The formalism offers potential for extension and generalization via multiple avenues. For instance, while our current treatment relies on a semiclassical description, therefore restricting it to a coherent state of light within the cavity, a valuable extension could involve capturing nonclassical photon statistics, e.g., by using cumulant expansion techniques Fowler-Wright et al. (2023). Other areas of interest are the study of spatially-resolved nonlinear polariton spectroscopy, which would involve generalizing the zero-dimensional cavity Eq. (3) to the 3D set of cavity modes afforded by classical electromagnetic simulations Taflove and Hagness (2005); Yang et al. (2023); Balasubrahmaniyam et al. (2023); Xu et al. (2023), or cavities exhibiting inherent nonlinearities Finkelstein-Shapiro et al. (2023). Furthermore, while we have mainly focused on pump-probe spectroscopy involving two pulses, the formalism can also be readily applied to 2D or multidimensional coherent spectroscopy by considering three or more pulses driving the cavity. Finally, let us emphasize that our approach is not limited to molecular polaritons, but could also be extended to other strongly coupled light-matter platforms, such as exciton-polaritons in inorganic semiconductor microcavities Fryett et al. (2018) or atomic gases.
Acknowledgments.—We are grateful to Wei Xiong for providing valuable feedback and comments on the manuscript. This research was primarily supported by the Air Force Office of Scientific Research (AFOSR) through the Multi-University Research Initiative (MURI) program no. FA9550-22-1-0317.